OpenStudy (anonymous):

2/(a+3) - (a^2+3a-10)/(a^3+5a^2+6a)=(a-4)/(a^2+3a) Solve for a. Please explaiiin to me how this is -6? And yes, the math answer key say's it's -6.

OpenStudy (anonymous):

I'm doing this but when I see other questions which don't involve as much time as this I jump to those. Don't worry, I'll have it in in a bit.

OpenStudy (anonymous):

2 / (a+3) - (a+5)(a-2) / (a+3)(a+2) = (a-4) / (a)(a+3) 2 - (a+5)(a-2) / (a+2) = (a-4) / a 2(a+2) / (a+2) - (a+5)(a-2) / (a+2) = (a-4) / a 2a + 4 - (a^2 +3a - 10) / (a+2) = (a-4) / a 2a + 4 - a^2 - 3a + 10 / (a+2) = (a-4) / a -a^2 - a + 14 / (a+2) = (a-4) / a -a^3 - a^2 + 14a = (a-4)(a+2) -a^3 - a^2 + 14a = a^2 - 2a - 8 -a^3 - 2a^2 + 14a + 8 = 0 a^3 + 2a^2 - 14a - 8 = 0 I don't think I'm getting -6. Probably did it too fast so I messed up somewhere.